# Instruction-Oriented Preference Alignment for Enhancing Multi-Modal Comprehension Capability of MLLMs

### v1

@@ Abstract
Preference alignment has emerged as an effective strategy to enhance the performance of Multimodal Large Language Models (MLLMs) following supervised fine-tuning. While existing preference alignment methods predominantly target hallucination factors, they overlook the factors essential for multi-modal comprehension capabilities, often narrowing their improvements on hallucination mitigation. To bridge this gap, we propose Instruction-oriented Preference Alignment (IPA), a scalable framework designed to automatically construct alignment preferences grounded in instruction fulfillment efficacy. Our method involves an automated preference construction coupled with a dedicated verification process that identifies instruction-oriented factors, avoiding significant variability in response representations. Additionally, IPA incorporates a progressive preference collection pipeline, further recalling challenging samples through model self-evolution and reference-guided refinement. Experiments conducted on Qwen2VL-7B demonstrate IPA’s effectiveness across multiple benchmarks, including hallucination evaluation, visual question answering, and text understanding tasks, highlighting its capability to enhance general comprehension.
@@ 3.4 Progressive Preference Data Collection
[…]
@@ 4.1 Implementation Details
Model Configurations. We adopt Qwen2VL-7B [68] as the base model architecture for preference data collection and performance validation, serving as the backbone for the generator, reviser, and verifier models. Parameters of different models are not shared. To enhance response diversity, we introduce additional models with comparable or weaker capabilities for response sampling (Equation 1-2) and feedback generation (Equation 3), including Qwen2VL-2B [68], InternVL2-2B&$\&8$[8], and LLaVA-OV-0.5B&\&&7B [28].
@@ 4.7 Generality of Our Method
For LLaVA-1.5-7B, we observe performance degradation when directly applying the curated preference data for alignment, which we conjecture is due to its constrained visual perception capabilities. Specifically, LLaVA-1.5-7B’s weaker vision encoder (limited to 336$\times$336 resolution) may struggle to percept the detailed visual elements in the data. Taking this into consideration, we simply filter out samples with response longer than 150 words, and create a subset (denoted as IPAsub) for the experiments.
@@ A.2 Prompt Design
We detail the prompts used in the automated preference construction. Specifically, these comprises the prompt $I^{fb}$ for feedback generation (Equation 3), the prompt $I^{rev}$ for response refinement (Equation 4), and the prompt $I^{ver}$ for instruction-oriented verification (Equation 5). The prompts are displayed in Table 5, 6 and 7 respectively.

### v2

@@ Abstract
Preference alignment has emerged as an effective strategy to enhance the performance of Multimodal Large Language Models (MLLMs) following supervised fine-tuning. While existing preference alignment methods predominantly target hallucination factors, they overlook the factors essential for multi-modal comprehension capabilities, often narrowing their improvements on hallucination mitigation. To bridge this gap, we propose Instruction-oriented Preference Alignment (IPA), a scalable framework designed to automatically construct alignment preferences grounded in instruction fulfillment efficacy. Our method involves an automated preference construction coupled with a dedicated verification process that identifies instruction-oriented factors, avoiding significant variability in response representations. Additionally, IPA incorporates a progressive preference collection pipeline, further recalling challenging samples through model self-evolution and reference-guided refinement. Experiments conducted on Qwen2VL-7B demonstrate IPA’s effectiveness across multiple benchmarks, including hallucination evaluation, visual question answering, and text understanding tasks, highlighting its capability to enhance general comprehension. The dataset is publicly available at https://huggingface.co/datasets/wangzt-kghl/IPA.
@@ 3.4 Progressive Preference Data Collection
Table 1: The evaluation results on different benchmarks. The upper part are the results reported in the original papers, while the lower part are our reproduced results (with $\dagger$}). Our method brings consistent improvements across 9 benchmarks compared to the base models, spanning three different dimensions: hallucination evaluation, general VQA, and text understanding.
@@ 4.1 Implementation Details
Model Configurations. We adopt Qwen2VL-7B [68] as the base model architecture for preference data collection and performance validation, serving as the backbone for the generator, reviser, and verifier models. Parameters of different models are not shared. To enhance response diversity, we introduce additional models with comparable or weaker capabilities for response sampling (Equation 1-2) and feedback generation (Equation 3), including Qwen2VL-2B [68], InternVL2-2B&8\&8[8], and LLaVA-OV-0.5B&\&7B [28].
@@ 4.7 Generality of Our Method
For LLaVA-1.5-7B, we observe performance degradation when directly applying the curated preference data for alignment, which we conjecture is due to its constrained visual perception capabilities. Specifically, LLaVA-1.5-7B’s weaker vision encoder (limited to $\times336$ resolution) may struggle to percept the detailed visual elements in the data. Taking this into consideration, we simply filter out samples with response longer than 150 words, and create a subset (denoted as IPAsub) for the experiments.
@@ A.2 Prompt Design
We detail the prompts used in the automated preference construction. Specifically, these comprises the prompt $I^{fb}$ for feedback generation (Equation 3), the prompt $I^{rev}$ for response refinement (Equation 4), and the prompt $I^{ver}$ for instruction-oriented verification (Equation 5). The prompts are displayed in Table 6, 7 and 8 respectively.
@@ Appendix B Qualitative Analysis
Appendix C More Experiments
C.1 Experiments with More Advanced Models
We also conduct experiments using larger and more advanced models, InternVL3-14B and Qwen2.5VL-72B, where the same dataset as in the main manuscript is used for training these models. As shown in Table 5, despite the preference dataset being constructed with a relatively weaker and smaller model (Qwen2VL-7B), it brings consistent gains to significantly larger models, demonstrating strong generalizability across model scales and architectures. While the improvements are more moderate due to the already strong baselines, our method still yields meaningful enhancements in overall comprehension ability. We believe that using stronger models for data collection can yield further improvements.